OpenStudy (anonymous):

I need to find the solution for these differential equations: 1) [(D^2)+D]y = x^2 + 3x + e^3x 2) y'' + 4y = 8cos2x -4x I know their complementary functions which are 1) yc = c1+c2e^-x 2) yc = c1cos2x +c2sin2x But I don't know how to get the particular functions to find the solution. Please help!!!

ganeshie8 (ganeshie8):

familiar with the method of undetermined coefficients ?

OpenStudy (anonymous):

yep. but i'm quite confused on how to use that

OpenStudy (michele_laino):

for the first one, try to search a solution like this: \[A{e^{3x}} + {B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3} + {B_5}{x^4}\]

ganeshie8 (ganeshie8):

can you use it on below two differential equations separately ? [(D^2)+D]y = x^2 + 3x [(D^2)+D]y = e^3x

OpenStudy (anonymous):

is this right? [(D^2)+D]y = e^3x y= c1+c2e^-x + (1/12)e^3x

OpenStudy (michele_laino):

yes!

ganeshie8 (ganeshie8):

yes only find a particular solution for each of the equations : [(D^2)+D]y = x^2 + 3x [(D^2)+D]y = e^3x

OpenStudy (anonymous):

but i don't know how to solve for this [(D^2)+D]y = x^2 + 3x

OpenStudy (michele_laino):

you have to consider a solution like this one: \[{B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3} + {B_5}{x^4}\]

OpenStudy (michele_laino):

we are applying the superposition principle of solutions

OpenStudy (anonymous):

i still can't understand ;;

OpenStudy (michele_laino):

you have to replace \(y'\) and \(y''\) with the first and second derivative of \({B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3} + {B_5}{x^4}\)

OpenStudy (michele_laino):

then you have to use the principle of identity of polynomials

OpenStudy (anonymous):

Is this right? B2=-1 B3=1/2 B4= 1/3 B5=0

ganeshie8 (ganeshie8):

you may use wolfram to double check http://www.wolframalpha.com/input/?i=solve+y%27%27%2By%27+%3D+x%5E2+%2B+3x+%2B+e%5E%283x%29 it seems you have got it correctly! good job !

OpenStudy (michele_laino):

correct!

OpenStudy (irishboy123):

the first one can be made less boring by multiplying by \(e^x\) and seeing the LHS as \( (e^x y^{\prime})^{\prime} \), n'est-ce pas?

OpenStudy (anonymous):

how about the B1?

OpenStudy (michele_laino):

\(B_1\) is added to \(c_1\) so we get the new constant: \(k_1=c_1+B_1\)

OpenStudy (anonymous):

thanks to all!!! can you still help me for the second one y'' + 4y = 8cos2x -4x??

OpenStudy (michele_laino):

for the second one, you have to try a solution like this one: \[{A_1}\cos \left( {2x} \right) + {A_2}\sin \left( {2x} \right) + {B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3}\]

OpenStudy (michele_laino):

or, if you prefer, you can split the problem, as @ganeshie8 has indicated before, and then try these solutions separately: \[{A_1}\cos \left( {2x} \right) + {A_2}\sin \left( {2x} \right)\] \[{B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3}\]

OpenStudy (michele_laino):

please wait, I have made a typo, the right solution is: \[x\left\{ {{A_1}\cos \left( {2x} \right) + {A_2}\sin \left( {2x} \right)} \right\} + {B_1} + {B_2}x + {B_3}{x^2} + {B_4}{x^3}\]

OpenStudy (anonymous):

is this the answer to the second one? y=c1cos2x+c2sin2x+2xsinx-x

OpenStudy (michele_laino):

do you mean: y=c1cos2x+c2sin2x+2xsin(2x)-x

OpenStudy (anonymous):

ah yes!

OpenStudy (michele_laino):

ok! correct!

OpenStudy (anonymous):

thanks a lot!!!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

can i still ask you a question?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Thanks. This is still a problem for differential equations. I have trouble solving these type of problems. Find the constant EMF E if C=(1/20)farad, R=20 ohms, q=0 when t=0, and q=2 coulomb when t=1 sec

OpenStudy (michele_laino):

is it a series RC circuit?

OpenStudy (anonymous):

ohm's law

OpenStudy (michele_laino):

we can write the subsequent ODE: \[ - \frac{Q}{C} + RI = {E_0}\cos \left( {\omega t} \right)\]

OpenStudy (michele_laino):

or, more precisely, we have this: \[\frac{{dV}}{{dt}} + \frac{V}{{RC}} = 0\] where \(V \) is the electric voltage

OpenStudy (michele_laino):

\(V= V(t)\) is the unknown function, of course

OpenStudy (michele_laino):

furthermore we have this equation: \[I = - C\frac{{dV}}{{dt}}\]

OpenStudy (michele_laino):

and, of course: \[Q\left( t \right) = CV\left( t \right)\]

OpenStudy (michele_laino):

those equations can solve your problem

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

:)